# Pace: Physics-Aware Attentive Temporal Convolutional Network for Battery Health Estimation

### v1

@@ 1. Introduction
Batteries are the foundations for modern energy systems, such as electric vehicles (EVs) and energy storage of the power grid. As electrification accelerates worldwide, battery demand is rising rapidly. According to the IEA Global EV Outlook 2025 (IEA, 2025), battery demand is largely driven by electric vehicles, e.g., lithium-ion batteries. The total demand for electric vehicles and energy storage reached 1 TWh in 2024, which is 25% more than in 2023 and is projected to triple to over 3 TWh by 2030. This surge in demand highlights the importance of robust battery health management, not just to prevent unexpected failures but also to extend battery life and improve system dependability and cost efficiency.
@@
Battery degrades over time, and the process is inevitable and complex. The degradation is influenced by many aspects, such as electrochemical reactions, charging/discharging patterns, and driving behaviors (Gao et al., 2022). Among various indicators of battery health, the state of health (SoH) is one of the most commonly used metrics and reflects the battery’s current capacity relative to its new battery nominal capacity (Chen et al., 2023). Accurate SoH estimation is vital for battery safety, lifespan, and efficiency. Despite growing interest, developing SoH prediction models that are accurate and robust under diverse operating conditions remains a non-trivial challenge. This is especially true when models need to be lightweight for real-world deployment in battery systems with limited computing resources.
@@
Traditional methods leverage domain knowledge to characterize battery behavior. For instance, the equivalent circuit model (ECM) represents a battery as a network of resistors and capacitors and is widely adopted for its simplicity and physical interpretability (Chen et al., 2023). Kalman filters are frequently applied to sensor measurements to estimate model parameters in real time (Zhao et al., 2025). More detailed physics-based approaches, such as the pseudo–two-dimensional model (Gao et al., 2022), replicate lithium-ion transport and provide insights into mechanisms like solid-electrolyte interphase (SEI) growth and lithium plating. While these methods offer valuable physical intuition, they often rely on expert tuning, handcrafted features, and simplified electrochemical approximations, which can limit scalability and adaptability under diverse operating conditions.
@@
Recently, data-driven and machine learning (ML) models have been proposed to learn complex degradation patterns directly from data using various algorithms. To model long sequences of battery data, methods such as long short-term memory networks (LSTMs) (Ouyang et al., 2024; Eleftheriadis et al., 2024), Transformers, temporal convolutional networks (TCNs) (Liu et al., 2024; Li et al., 2023), and graph neural networks (Yao et al., 2023) have achieved competitive performance in battery analytics with their ability to capture long, nonlinear, and inconsistent behaviors of battery degradation. Hybrid models have also emerged to combine complementary strengths (Sameer et al., 2025). For example, incremental capacity analysis (Xu et al., 2024) derives differential capacity curves from voltage–capacity data and adds them to ML feature spaces to improve accuracy. Physics-informed neural networks (Wang et al., 2024) have also been used to include physical constraints or governing equations into ML architectures or loss functions to improve prediction robustness, reduce data dependence, and enforce physically consistent behavior. However, because battery aging involves multiple coupled and nonlinear processes, formulating full partial differential equation models with tractable boundary conditions remains challenging in practice. In contrast, simpler empirical models like the ECM offer interpretable parameters without requiring full electrochemical modeling.
@@ 2. Battery Physics Modelling with ECM
A key strength of Pace is its usage of battery physics to complement raw sensor data. While common battery measurements such as voltage and current offer snapshots of battery dynamics, they lack the ability to capture internal battery electrochemical dynamics, which evolve with battery degradation. To better capture the battery’s transient behavior, Pace extracts physics features from ECM and utilizes the features, compact yet expressive, for SoH prediction. Among various ECMs, we adopt the widely used first-order Thevenin model (Tran et al., 2021) for its well-balanced trade-off between accuracy and efficiency. The model represents the battery as a combination of an ideal voltage source, internal resistance, and capacitance to reveal transient behavior with high modelling accuracy. The efficiency of the model allows us to inject physically meaningful features into Pace to achieve improved robustness and generalization without incurring significant computational overhead. Specifically, the first-order model consists of a few key components that collectively represent the battery’s electrical behavior under various operating conditions (Tao et al., 2024). The first is the OCV, which is the battery’s ideal voltage when no current is drawn. Next is the Ohmic resistance $R_{0}$, which models voltage’s immediate drop when current is drawn, i.e., the battery’s instantaneous response. The slow and transient changes in voltage are captured by a resistor-capacitor (RC) pair. It simulates the internal resistive and capacitive delays that prevent the battery from reaching expected voltage during charging or discharging. The model dynamics are described by the below equations.
@@ 3.1. Dilated Temporal Block (DTB)
Battery degradation is progressive and non-linear over time. To effectively learn the temporal dependencies embedded in both the ECM-derived features and raw data, we place the DTB immediately after the input to capture long-range patterns early in the pipeline of Pace, as shown in Figure 2(B). DTB is used for sequential modelling based on dilated causal convolutions (Terzic et al., 2023; Lai et al., 2023). We follow the general structure of DTB and introduce battery-specific customizations detailed as follows.
@@ 3.2. Chunked Attention Block (CAB)
DTB alone, based on local convolution, cannot well capture gradual and non-local patterns over the battery degradation horizons. While it captures structured dependencies through expanding receptive fields, it lacks the flexibility to adjust the weights or focus on different parts of the history data as degradation evolves. This motivates the use of a scalable attention mechanism to incorporate additional context. Full attention is often the default choice; however, it can be computationally expensive for long time-series sequences with its quadratic complexity $\mathcal{O}(n^{2})$ with sequence length $n$. To incorporate scalable attention within temporal modeling, we introduced CAB (Zeineldeen et al., 2024) into the pipeline, as presented in Figure 2(C).
@@ 3.3. Multi-Scale Dual-Head Output Block (DHB)
Degradation evolves nonlinearly across a battery’s lifecycle. It often starts slowly for new batteries and accelerates towards the batteries’ EoL. Among different batteries, the capacity decay may show sharp drops, plateaus, or irregularities due to thermal, load, or manufacturing variations. Accurately predicting SoH thus requires the model to comprehensively consider degradations in different stages and battery settings. This should be addressed not only in hidden layers, e.g., in DTB and CAB, but also in output layers. In many predictive models, only a single linear decoder is configured in the output layer to map the extracted features from the hidden layers to the output. The simple design often struggles to generalize across different degradation patterns, e.g., over-smooth rapid decay and overfit to local noises. To address the challenges, we adopt a multi-head concept (Zhou et al., 2021) and propose DHB with two complementary heads, as depicted in Figure 2(D). The first head is convolutional-based. It applies a 1D convolution over a fixed-size window to capture local and fine-grained temporal SoH dynamics. It also has a linear head, which configures a simple feedforward layer to capture global or smoothed trends and map the latent state to SoH prediction. The outputs of the two heads are fused into one, and the weight of each output is controlled by a learnable scalar gate $\alpha\in[0,1]$. Let $\hat{y}_{\text{conv}}$ and $\hat{y}_{\text{linear}}$ be the outputs of the convolutional and linear heads, respectively. The predicted SoH is computed as,
@@ 4.1.1. Dataset
We choose one of the largest public battery datasets from (Severson et al., 2019) for our study. The dataset consists of 124 commercial Li-ion battery cells cycled under various charging/discharging conditions with diverse degradation patterns. Each cell has a nominal capacity of 1.1 Ah and was tested in a temperature-controlled chamber (30°C) until it reached 80% of its initial capacity, i.e., EoL. The lifetime of the batteries varies between 150 and 2,300 cycles. Each cycle includes time-series sensor measurements for voltage, current, and temperature. SoH at cycle $t$ is given by $y_{\text{SoH}}^{t}=(Q_{t}/Q_{0})\times 100\%$, where $Q_{t}$ and $Q_{0}$ are the maximum available capacity at cycle $t$ and initial capacity before cycling. SoH is modeled as a percentage, which shall not exceed 100%, e.g., optimal condition for new batteries. To ensure a realistic evaluation, we follow the latest settings of (Xie et al., 2025), where 42 cells across 18 distinct charging protocols are carefully selected and used for training, and 6 cells with different charging conditions are selected for testing. Such a controlled configuration of training and testing data enables consistent and fair performance comparison.
@@ 4.1.3. Metrics
Pace is designed with real-world deployment in mind. Instead of focusing solely on accuracy, as many academic studies do, we follow applied settings and evaluate both prediction accuracy and model efficiency, which is a key requirement for edge devices in battery health monitoring. Specifically, we not only adopt commonly used accuracy metrics, including mean absolute error (MAE) and root mean squared error (RMSE), but also quantify efficiency through model size and FLOPs. In addition, we are inspired by a recent work (Zhao et al., 2023) and define a metric $\eta$ called efficiency as,
@@ 4.2.1. Comparison Algorithms
We compare Pace with several representative baselines, each of which reflects a different modeling philosophy. First, we choose a model that is accuracy-driven with no restrictions on model size. We expect our proposed Pace to approximate the accuracy of this model as much as possible and, meanwhile, maintain its advantage in practical edge deployment. Specifically, we choose Transformer (Sun et al., 2025), which is attention-based and offers high prediction accuracy for SoH. However, it has significant computational overhead and is impractical for embedded battery systems. Second, we prioritize model size and choose TCN. It is a lightweight model based on dilated causal convolutions and known for its efficiency. We expect Pace to be as compact as possible in comparison with TCN and able to perform well for both short-term and long-term SoH prediction. Finally, we choose one of the latest models, BiLSTM-CNN-Attention (Xie et al., 2025), referred to as BCA for simplicity. The model has a hybrid design and combines recurrent, convolutional, and attention mechanisms. It is one of the most competitive models for SoH prediction based on our survey.

### v2

@@ 1. Introduction
Batteries are the foundations for modern energy systems, such as electric vehicles (EVs) and energy storage of the power grid. As electrification accelerates worldwide, battery demand is rising rapidly. According to the IEA Global EV Outlook 2025 (IEA25), battery demand is largely driven by electric vehicles, e.g., lithium-ion batteries. The total demand for electric vehicles and energy storage reached 1 TWh in 2024, which is 25% more than in 2023 and is projected to triple to over 3 TWh by 2030. This surge in demand highlights the importance of robust battery health management, not just to prevent unexpected failures but also to extend battery life and improve system dependability and cost efficiency.
@@
Battery degrades over time, and the process is inevitable and complex. The degradation is influenced by many aspects, such as electrochemical reactions, charging/discharging patterns, and driving behaviors (9384220). Among various indicators of battery health, the state of health (SoH) is one of the most commonly used metrics and reflects the battery’s current capacity relative to its new battery nominal capacity (chen2023state). Accurate SoH estimation is vital for battery safety, lifespan, and efficiency. Despite growing interest, developing SoH prediction models that are accurate and robust under diverse operating conditions remains a non-trivial challenge. This is especially true when models need to be lightweight for real-world deployment in battery systems with limited computing resources.
@@
Traditional methods leverage domain knowledge to characterize battery behavior. For instance, the equivalent circuit model (ECM) represents a battery as a network of resistors and capacitors and is widely adopted for its simplicity and physical interpretability (chen2023state). Kalman filters are frequently applied to sensor measurements to estimate model parameters in real time (11065210). More detailed physics-based approaches, such as the pseudo–two-dimensional model (9384220), replicate lithium-ion transport and provide insights into mechanisms like solid-electrolyte interphase (SEI) growth and lithium plating. While these methods offer valuable physical intuition, they often rely on expert tuning, handcrafted features, and simplified electrochemical approximations, which can limit scalability and adaptability under diverse operating conditions.
@@
Recently, data-driven and machine learning (ML) models have been proposed to learn complex degradation patterns directly from data using various algorithms. To model long sequences of battery data, methods such as long short-term memory networks (LSTMs) (ouyang2024combined; eleftheriadis2024joint), Transformers, temporal convolutional networks (TCNs) (liu2024state; li2023makes), and graph neural networks (yao2023novel) have achieved competitive performance in battery analytics with their ability to capture long, nonlinear, and inconsistent behaviors of battery degradation. Hybrid models have also emerged to combine complementary strengths (sameer2025ginet). For example, incremental capacity analysis (xu2024state) derives differential capacity curves from voltage–capacity data and adds them to ML feature spaces to improve accuracy. Physics-informed neural networks (wang2024physics) have also been used to include physical constraints or governing equations into ML architectures or loss functions to improve prediction robustness, reduce data dependence, and enforce physically consistent behavior. However, because battery aging involves multiple coupled and nonlinear processes, formulating full partial differential equation models with tractable boundary conditions remains challenging in practice. In contrast, simpler empirical models like the ECM offer interpretable parameters without requiring full electrochemical modeling.
@@ 2. Battery Physics Modelling with ECM
A key strength of Pace is its usage of battery physics to complement raw sensor data. While common battery measurements such as voltage and current offer snapshots of battery dynamics, they lack the ability to capture internal battery electrochemical dynamics, which evolve with battery degradation. To better capture the battery’s transient behavior, Pace extracts physics features from ECM and utilizes the features, compact yet expressive, for SoH prediction. Among various ECMs, we adopt the widely used first-order Thevenin model (tran2021comparative) for its well-balanced trade-off between accuracy and efficiency. The model represents the battery as a combination of an ideal voltage source, internal resistance, and capacitance to reveal transient behavior with high modelling accuracy. The efficiency of the model allows us to inject physically meaningful features into Pace to achieve improved robustness and generalization without incurring significant computational overhead. Specifically, the first-order model consists of a few key components that collectively represent the battery’s electrical behavior under various operating conditions (tao2024state). The first is the OCV, which is the battery’s ideal voltage when no current is drawn. Next is the Ohmic resistance $R_{0}$, which models voltage’s immediate drop when current is drawn, i.e., the battery’s instantaneous response. The slow and transient changes in voltage are captured by a resistor-capacitor (RC) pair. It simulates the internal resistive and capacitive delays that prevent the battery from reaching expected voltage during charging or discharging. The model dynamics are described by the below equations.
@@ 3.1. Dilated Temporal Block (DTB)
Battery degradation is progressive and non-linear over time. To effectively learn the temporal dependencies embedded in both the ECM-derived features and raw data, we place the DTB immediately after the input to capture long-range patterns early in the pipeline of Pace, as shown in Figure 2(B). DTB is used for sequential modelling based on dilated causal convolutions (terzic2023tcnca; lai2023lightcts). We follow the general structure of DTB and introduce battery-specific customizations detailed as follows.
@@ 3.2. Chunked Attention Block (CAB)
DTB alone, based on local convolution, cannot well capture gradual and non-local patterns over the battery degradation horizons. While it captures structured dependencies through expanding receptive fields, it lacks the flexibility to adjust the weights or focus on different parts of the history data as degradation evolves. This motivates the use of a scalable attention mechanism to incorporate additional context. Full attention is often the default choice; however, it can be computationally expensive for long time-series sequences with its quadratic complexity $\mathcal{O}(n^{2})$ with sequence length $n$. To incorporate scalable attention within temporal modeling, we introduced CAB (zeineldeen2024chunked) into the pipeline, as presented in Figure 2(C).
@@ 3.3. Multi-Scale Dual-Head Output Block (DHB)
Degradation evolves nonlinearly across a battery’s lifecycle. It often starts slowly for new batteries and accelerates towards the batteries’ EoL. Among different batteries, the capacity decay may show sharp drops, plateaus, or irregularities due to thermal, load, or manufacturing variations. Accurately predicting SoH thus requires the model to comprehensively consider degradations in different stages and battery settings. This should be addressed not only in hidden layers, e.g., in DTB and CAB, but also in output layers. In many predictive models, only a single linear decoder is configured in the output layer to map the extracted features from the hidden layers to the output. The simple design often struggles to generalize across different degradation patterns, e.g., over-smooth rapid decay and overfit to local noises. To address the challenges, we adopt a multi-head concept (zhou2021informer) and propose DHB with two complementary heads, as depicted in Figure 2(D). The first head is convolutional-based. It applies a 1D convolution over a fixed-size window to capture local and fine-grained temporal SoH dynamics. It also has a linear head, which configures a simple feedforward layer to capture global or smoothed trends and map the latent state to SoH prediction. The outputs of the two heads are fused into one, and the weight of each output is controlled by a learnable scalar gate $\alpha\in[0,1]$. Let $\hat{y}_{\text{conv}}$ and $\hat{y}_{\text{linear}}$ be the outputs of the convolutional and linear heads, respectively. The predicted SoH is computed as,
@@ 4.1.1. Dataset
We choose one of the largest public battery datasets from (severson2019data) for our study. The dataset consists of 124 commercial Li-ion battery cells cycled under various charging/discharging conditions with diverse degradation patterns. Each cell has a nominal capacity of 1.1 Ah and was tested in a temperature-controlled chamber (30°C) until it reached 80% of its initial capacity, i.e., EoL. The lifetime of the batteries varies between 150 and 2,300 cycles. Each cycle includes time-series sensor measurements for voltage, current, and temperature. SoH at cycle $t$ is given by $y_{\text{SoH}}^{t}=(Q_{t}/Q_{0})\times 100\%$, where $Q_{t}$ and $Q_{0}$ are the maximum available capacity at cycle $t$ and initial capacity before cycling. SoH is modeled as a percentage, which shall not exceed 100%, e.g., optimal condition for new batteries. To ensure a realistic evaluation, we follow the latest settings of (xie2025data), where 42 cells across 18 distinct charging protocols are carefully selected and used for training, and 6 cells with different charging conditions are selected for testing. Such a controlled configuration of training and testing data enables consistent and fair performance comparison.
@@ 4.1.3. Metrics
Pace is designed with real-world deployment in mind. Instead of focusing solely on accuracy, as many academic studies do, we follow applied settings and evaluate both prediction accuracy and model efficiency, which is a key requirement for edge devices in battery health monitoring. Specifically, we not only adopt commonly used accuracy metrics, including mean absolute error (MAE) and root mean squared error (RMSE), but also quantify efficiency through model size and FLOPs. In addition, we are inspired by a recent work (zhao2023deep) and define a metric $\eta$ called efficiency as,
@@ 4.2.1. Comparison Algorithms
We compare Pace with several representative baselines, each of which reflects a different modeling philosophy. First, we choose a model that is accuracy-driven with no restrictions on model size. We expect our proposed Pace to approximate the accuracy of this model as much as possible and, meanwhile, maintain its advantage in practical edge deployment. Specifically, we choose Transformer (sun2025optimizing), which is attention-based and offers high prediction accuracy for SoH. However, it has significant computational overhead and is impractical for embedded battery systems. Second, we prioritize model size and choose TCN. It is a lightweight model based on dilated causal convolutions and known for its efficiency. We expect Pace to be as compact as possible in comparison with TCN and able to perform well for both short-term and long-term SoH prediction. Finally, we choose one of the latest models, BiLSTM-CNN-Attention (xie2025data), referred to as BCA for simplicity. The model has a hybrid design and combines recurrent, convolutional, and attention mechanisms. It is one of the most competitive models for SoH prediction based on our survey.